OpenStudy (anonymous):

converge? diverge? \[\frac{ 1 }{ 2^n -n }\]

OpenStudy (jhannybean):

n goes to infinity?

OpenStudy (anonymous):

oops sorry. n goes to 1

OpenStudy (jhannybean):

Can you write out the series? It's a little confusing...:\

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ 2^n -n }\]

OpenStudy (anonymous):

here it is

OpenStudy (jhannybean):

Yay

OpenStudy (jhannybean):

i'm thinking you can use a Direct comparison test here, to compare \(\large a_{n}\) to \(\large b_{n}\)

OpenStudy (jhannybean):

\[\large a_{n} =\frac{1}{2^{n}-n} \ , \ b_{n}= \frac{1}{2^{n}}\] \(\large b_{n}\) converges because \(\large\frac12 < 1\) Direct comparison test states that if \(\large a_{n} < b_{n} \ \text{AND} \ \ \sum{b_{n}} = C \) then \(\large \sum{a_{n}} = c\) We can test that \(\large \frac{1}{2^{n}-n} < \frac{1}{2^{n}}\) therefore \[\large \sum_{n=1}^{\infty} \frac{1}{2^n -n} =C\] by Direct Comparison Test.